OpenStudy (anonymous):

Factor: x^2 − 4x + 5

jhonyy9 (jhonyy9):

x_1,_2=(4+/-sqrt(16-20))/2=(4+/-sqrt(-4))/2=(4+/-(2i))/2=2+/-i x_1=2-i x_2=2+i (x-(2-i))(x-(2+i)) (x-2+i)(x-2-i)